# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v1

@@ A.2.1 Unit Root Regressors
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi^{(j)*})\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$
Therefore, (A.34) holds as $\mu=C_{\mu}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ with a sufficiently large $C_{\mu}$. This completes the proof of Proposition 1. ∎
@@ A.2.2 Stationary Regressor
Note that $\phi_{j,t-1}\in\sigma(w_{j,0},\dots,w_{j,t-2}).$ Then $\phi_{j,t}$ is an AR(1) process with coefficient $1-C_{\zeta}/n^{\tau}$ and innovation $w_{j,t-1}$. Recall that $w_{j,t}$ is $\alpha$-mixing and sub-exponential; we obtain (A.46) following the same arguments for (A.14),
@@
This completes the proof of Proposition 2. ∎
@@
Note that in (A.36), we allow a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(2C_{\sigma}n^{1-\tau})$ with the absolute constant $C_{\sigma}$ in (A.36). Then
@@
$\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{2C_{\sigma}C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
@@
$\|\sum_{t=1}^{n}D_{-j}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\leq\dfrac{C_{\mu}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$ (A.81)
@@ Appendix B Additional Empirical Results
AR(1) Coef. ADF $p$-value AR(1) Coef. ADF $p$-value
Full Sample 0.224 $<$0.01 0.993 0.081
(Jan. 1960 - Dec. 2019)
Pre-1994 0.255 $<$0.01 0.994 0.776
(Jan. 1960 - Dec. 1993)
Post-1994 0.193 $<$0.01 0.979 $<$0.01
(Jan. 1994 - Apr. 2024)
@@
AR(1) Coef. ADF $p$-value AR(1) Coef. ADF $p$-value
Full Sample 0.621 0.021 1.000 0.081
(Jan. 1960 - Dec. 2019)
Pre-Volcker 0.640 0.220 0.986 0.406
(Jan. 1960 - July 1979)
Volcker-Greenspan 0.623 $<$0.01 0.992 0.071
(Aug. 1979 - Jan. 2006)
Bernanke-Yellen-Powell 0.543 $<$0.01 0.938 0.365
(Feb. 2006 - Apr. 2024)

### v2

@@ A.2.1 Unit Root Regressors
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$
Therefore, (A.34) holds as $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ with a sufficiently large $C_{{\rm a}}$. This completes the proof of Proposition 1 for $j\in\mathcal{M}_{x}$. ∎
@@ A.2.2 Stationary Regressor
Note that $\phi_{j,t-1}\in\sigma(w_{j,0},\dots,w_{j,t-2}).$ Then $\phi_{j,t}$ is an AR(1) process with coefficient $1-C_{\zeta}/n^{\tau}$ and innovation $w_{j,t-1}$. Recall that $w_{j,t}$ is $\alpha$-mixing and sub-exponential; we obtain (A.47) following the same arguments for (A.14).
@@
This completes the proof of Proposition 1 for $j\in\mathcal{M}_{z}$. ∎
@@
Note that in (A.36), we allow for a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(4C_{\sigma}s^{2}n^{1-\tau})$ with the absolute constant $C_{\sigma}$ in (A.36). Then
@@
$\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
@@
$\|\sum_{t=1}^{n}D_{-j}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$ (A.82)
@@ Appendix B Additional Empirical Results
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2024) 0.224 $<$0.01 0.993 0.081
Pre-1994 (Jan. 1960 - Dec. 1993) 0.255 $<$0.01 0.994 0.776
Post-1994 (Jan. 1994 - Apr. 2024) 0.193 $<$0.01 0.979 $<$0.01
@@
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2024) 0.621 0.021 1.000 0.081
Pre-Volcker (Jan. 1960 - July 1979) 0.640 0.220 0.986 0.406
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.623 $<$0.01 0.992 0.071
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2024) 0.543 $<$0.01 0.938 0.365